OpenStudy (anonymous):

What value of n solves the equation? 3^n=1/81 A. –243 B. –27 C. –4 D. 4

OpenStudy (anonymous):

@WisdomDaughter @guillermo2cruzIII

OpenStudy (gorv):

\[3^{n}=\frac{ 1 }{ 81 }=\frac{ 1 }{ 3^4 }=3^{-4}\]

OpenStudy (gorv):

both side have same base=3 these can be equal only when their power will be equal

OpenStudy (gorv):

so n=-4

OpenStudy (anonymous):

Thank you thank you thank you!!!! You just saved me from failing a math test!!!! @gorv